Problem Page
In the lab, Heather has two solutions that contain alcohol and is mixing them with each other. She uses
500
milliliters less of Solution A than Solution B. Solution A is
17%
alcohol and Solution B is
11%
alcohol. How many milliliters of Solution A does she use, if the resulting mixture has
167
milliliters of pure alcohol?

Found 2 solutions by ankor@dixie-net.com, stanbon:
Answer by ankor@dixie-net.com(22740)

In the lab, Heather has two solutions that contain alcohol and is mixing them with each other.
She uses 500 milliliters less of Solution A than Solution B.
Solution A is 17% alcohol and
Solution B is 11% alcohol.
How many milliliters of Solution A does she use, if the resulting mixture has
167 milliliters of pure alcohol?
:
a = amt of A solution
b = amt of B solution
"500 milliliters less of Solution A than Solution B."
a = b - 500
:
A mixture equation
.17a + .11b = 167
Replace a with (b-500)
.17(b-500) + .11b = 167
.17b - 85 + .11b = 167
.17b + .11b = 167 + 85
.28b = 252
b = 252/,28
b = 900 ml of B solution used
therefore
900 - 500 = 400 ml of A solution
:
;
Check by finding the actual amt of alcohol in each
.17(400) = 68 ml
.11(900) = 99
------------------------
total alcohol: 167 mo

Answer by stanbon(75887)

Heather has two solutions that contain alcohol and is mixing them with each other. She uses 500 milliliters less of Solution A than Solution B.
Solution A is 17% alcohol and Solution B is 11% alcohol.
How many milliliters of Solution A does she use, if the resulting mixture has
167 milliliters of pure alcohol?
------------
B - A = 500
0.11B + 0.17A = 1*167
----------------------
11B - 11A = 11*500
11B + 17A = 100*167
-----------------
6A = 22200
A = 3700 ml
------------
Cheers,
Stan H.
------------